OpenStudy (anonymous):

Evaluate the expression below when x = -2. X^2-8/x+6

OpenStudy (anonymous):

\[x^{2}-8/x+6\]

jigglypuff314 (jigglypuff314):

plug in (-2) for x so (-2)^2 - 8(-2) + 6 = ?

OpenStudy (anonymous):

May you help me with that

OpenStudy (anonymous):

The possible answers are -3 -1 1/2 3/2

OpenStudy (liamschumm):

Here is a hint. \[(-2)^2 - 8(-2) + 6=2^2 +8*2 + 6.\]

OpenStudy (liamschumm):

Wait, no.

OpenStudy (liamschumm):

(-2)^2 - 8/(-2) + 6.

OpenStudy (liamschumm):

2^2 + 8/2 + 6 = ?